Faith mixes 14 liters of 12% acid solution with a 30% acid solution to make an 18% acid solution. How many liters of 30% solution did she use?
Answer by TimothyLamb(4379)

x = liters of 30% acid
---
0.12*14 + 0.30x = 0.18(x + 14)
1.68 + 0.30x = 0.18x + 2.52
0.30x - 0.18x = 2.52 - 1.68
0.12x = 2.52 - 1.68
x = (2.52 - 1.68)/0.12
x = 7
---
answer:
7 liters of 30% acid
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---